# The Elephant in the Room: Rethinking the Usage of Pre-trained Language Model in Sequential Recommendation

### v2

@@ 3.3. Analysis on Model Redundancy
Conventional ID-based SR models typically feature a simplified architecture with fewer layers and heads (for example, SASRec employs 2 layers and 1 head). As reported in their respective studies (SASRec; BERT4Rec), augmenting the number of attention heads and layers tends to result in a marginal performance improvement and, in some cases, can even cause a reduction in effectiveness. However, PLMs usually have several times more layers and heads (e.g., Longformer has 12 layers and 12 heads), aiming to capture the complex grammar and semantic knowledge within natural language. This level of complexity might be over-qualified for the current requirements of behavior sequence modeling in SR. The observed attention homogenization on different layers and heads also indicates that the current sequential modeling of PLM-based SR models is redundant and could be compressed. Hence, we conduct an intuitive experiment that only tunes several layers of RECFORMER on Amazon Pantry as shown in Table 1. We have some findings:
@@ 4.1. Behavior-tuned PLMs as Item Initializer
We conduct extensive experiments on the SOTA PLM-based SR model RECFORMER (li2023text), and implement our enhanced variants with simplified sequence modeling based on it. We adopt two classical SR models SASRec (SASRec) and BERT4Rec (BERT4Rec) as the backbone sequence modeling method. Our PLM-based SR variants can be divided into three groups, namely embedding-init (freeze), embedding-init (trainable), and further-train. Note that we use the behavior-tuned PLM (i.e., Longformer tuned in RECFORMER’s certain stages in Sec. 3.1) to initialize item embeddings in our methods.
(a) For embedding-init (freeze), we experiment with three versions of RECFORMER: the original Longformer (longformer) (LF), RECFORMER trained on all pre-training datasets (PT), and RECFORMER after stage-FT1 on the target domain (FT). We only use the item representations obtained by these models as the fixed item ID embeddings of ID-based SR models (e.g., SASRec), and fine-tune the sequence modeling part of ID-based SR models via the target domain’s training data. This setting could be viewed as an SR model with fixed item embeddings given by Longformer/RECFORMER, aiming to explore whether a simple sequence model from conventional ID-based SR is sufficient for behavior modeling of SR.
@@ 4.2. Experimental Settings
Datasets. We conduct experiments on seven real-world datasets from the Amazon review datasets  (amazon). We use Books, Movies and TV, Sports and Outdoors, Clothing Shoes and Jewelry for pre-training, and choose Arts, Crafts and Sewing, Musical Instruments, and Pantry as the downstream domains for evaluation. In addition, we also select Movies and TV and Sports and Outdoors from the pre-training dataset for evaluation to test the impact of pre-training on the source domain. Following (SASRec; BERT4Rec), for pre-training datasets, we filter out users and items with less $5$ interactions. But for new domain datasets, we just filter out the users with less than 4 interaction records to ensure adequate instances. We arrange each user’s historical behavior sequence by timestamps and adopt leave-one-out setting for building validation and test sets. For each user historical sequence, the last item is used as the test data, the item before the last one is used as the validation data, and the remaining interaction sequences are used for training.
@@
Evaluation Metrics. To evaluate the performance of the Sequential Recommendation task, we follow the classical setting (SASRec; BERT4Rec) and select $100$ randomly sampled negative items for evaluation. We adopt two widely-used metrics Hit rate (HR@k) and Normalized Discounted Cumulative Gain (NDCG@k) with $k=5,10$.
@@ 4.3. Main Results
(5) Sequence modeling methods that possess analogous architectures and training objectives to those of behavior-tuned PLMs benefit more from our PLM-based initialization. Our item initialization is universal for different ID-based sequence modeling. Comparing the results based on SASRec and BERT4Rec, we find that our initialization facilitates more for BERT4Rec under the freeze setting. We hypothesize that it is attributed to the similarities in sequence modeling architectures and training losses between RECFORMER and BERT4Rec. It also supports our hypothesis that PLMs may be excessively complex for behavior sequence modeling and could be substituted with simpler sequence models.
@@ 4.4. Results of Full-ranking Setting
The findings in Sec. 4.3 demonstrate the effectiveness of our proposed behavior-tuned PLMs initialized item embeddings under the random negative sampling setting. To further assess the robustness of our method, we conduct experiments with the full-ranking setting (metric; metric2) on several datasets. From Table 4, we have:
@@
(2) Our method still consistently outperforms RECFORMER on all datasets. Compared with the random-sampled setting in Table 3, the improvement in the full-ranking setting is not that impressive. This phenomenon is consistent with previous works (li2023text; qu2023thoroughly) that PLMs-based SR models perform well at making precise recommendations (i.e., pushing positive samples to the very top part of all candidates). It should be highlighted that our method greatly reduces the training and inference costs compared with RECFORMER. We only use PLM for item initialization, and the subsequent tuning and inference are only based on simplified sequence models without PLM. Considering our simple structure in serving (the same as SASRec), the current improvement is acceptable. The advantage of our method might be more activated in the matching stage (selecting the top hundreds of items from millions of candidates).
@@ 4.5. Analysis on Further Cooperations
Cooperating with trainable ID embeddings. Some previous works enhance item representation by adding the textual embeddings to its trainable item ID embeddings (UniRec; IDA-SR). Following this, we regard our behavior-tuned PLM initialization as the extra textual embeddings in SASRec and fix it during training (i.e., initialized embeddings as additional training signals). The results in Table 5 demonstrate that employing behavior-tuned PLMs initialized embeddings (i.e., PT and FT) leads to substantial performance enhancements, while embeddings initialized by original PLMs (LF) do not. The results reconfirm that it is the behavior-aware semantics, not the original semantic information, that improves the SR performance.
@@
4.6. Analysis on Pre-training Sizes.
The Pre-training datasets have been verified to be beneficial for RECFORMER in SR tasks on new domains. In Table 3, our PT setting also achieves satisfactory results. In this section, we explore the influence of pre-training dataset size on the effectiveness of our initialization by conducting experiments with PLMs pre-trained on randomly selected subsets of 0.01%, 0.1%, 1%, 2%, 10%, 20%, and 50% from the original pre-training dataset. From Fig. 2, we can observe that leveraging pre-training on behavioral datasets could significantly improve our performance. It is also astonishing that even a small amount of behavior-based pre-training (e.g., involving several thousand users) could successfully add the behavior-level preference to PLMs and achieve large improvement compared to the original Longformer. This observation reinforces the conclusion that the effectiveness of our method is primarily attributed to behavior-aware PLMs rather than the original PLMs. It also suggests that a minimal amount of pre-training behavioral data is possibly sufficient and transferable for new-domain SR tasks, implying that our method is practical and efficient based on a unified behavior-tuned PLM-based model.
@@ 5. CONCLUSION and FUTURE WORK
In this work, we deeply explore the effect of PLMs in SR, identifying significant model underutilization and redundancy of PLM in behavior sequence modeling by attention analyses and freezing/simplifying the sequence modeling of PLM. Next, we conduct extensive experiments and analyses on different PLM-based SR variants, discovering that the simplified sequence modeling of conventional ID-based SR models enhanced with the proposed behavior-tuned PLMs initialized item embeddings achieves a giant performance boost, while vanilla PLMs initializations are unable to obtain any improvements.
In the future, we will explore the scalability of both SR datasets (including the size and domain diversity) and PLMs, determining whether any scaling laws apply to our proposed initialization. Additionally, identifying more appropriate training objectives to align PLMs with behavioral knowledge for initialization presents a promising avenue for research.
[…]

### v3

@@ 3.3. Analysis on Model Redundancy
Conventional ID-based SR models typically feature a simplified architecture with fewer layers and heads (for example, SASRec employs 2 layers and 1 head). As reported in their respective studies (Kang and McAuley, 2018; Sun et al., 2019), augmenting the number of attention heads and layers tends to result in a marginal performance improvement and, in some cases, can even cause a reduction in effectiveness. However, PLMs usually have several times more layers and heads (e.g., Longformer has 12 layers and 12 heads), aiming to capture the complex grammar and semantic knowledge within natural language. This level of complexity might be over-qualified for the current requirements of behavior sequence modeling in SR. The observed attention homogenization on different layers and heads also indicates that the current sequential modeling of PLM-based SR models is redundant and could be compressed. Hence, we conduct an intuitive experiment that only tunes several layers of RECFORMER on Amazon Pantry as shown in Table 1. We have some findings:
@@ 4.1. Behavior-tuned PLMs as Item Initializer
We conduct extensive experiments on the SOTA PLM-based SR model RECFORMER (Li et al., 2023), and implement our enhanced variants with simplified sequence modeling based on it. We adopt two classical SR models SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) as the backbone sequence modeling method. Our PLM-based SR variants can be divided into three groups, namely embedding-init (freeze), embedding-init (trainable), and further-train. Note that we use the behavior-tuned PLM (i.e., Longformer tuned in RECFORMER’s certain stages in Sec. 3.1) to initialize item embeddings in our methods.
(a) For embedding-init (freeze), we experiment with three versions of RECFORMER: the original Longformer (Beltagy et al., 2020) (LF), RECFORMER trained on all pre-training datasets (PT), and RECFORMER after stage-FT1 on the target domain (FT). We only use the item representations obtained by these models as the fixed item ID embeddings of ID-based SR models (e.g., SASRec), and fine-tune the sequence modeling part of ID-based SR models via the target domain’s training data. This setting could be viewed as an SR model with fixed item embeddings given by Longformer/RECFORMER, aiming to explore whether a simple sequence model from conventional ID-based SR is sufficient for behavior modeling of SR.
@@ 4.2. Experimental Settings
Datasets. We conduct experiments on seven real-world datasets from the Amazon review datasets  (Ni et al., 2019). We use Books, Movies and TV, Sports and Outdoors, Clothing Shoes and Jewelry for pre-training, and choose Arts, Crafts and Sewing, Musical Instruments, and Pantry as the downstream domains for evaluation. In addition, we also select Movies and TV and Sports and Outdoors from the pre-training dataset for evaluation to test the impact of pre-training on the source domain. Following (Kang and McAuley, 2018; Sun et al., 2019), for pre-training datasets, we filter out users and items with less $5$ interactions. But for new domain datasets, we just filter out the users with less than 4 interaction records to ensure adequate instances. We arrange each user’s historical behavior sequence by timestamps and adopt leave-one-out setting for building validation and test sets. For each user historical sequence, the last item is used as the test data, the item before the last one is used as the validation data, and the remaining interaction sequences are used for training.
@@
Evaluation Metrics. To evaluate the performance of the Sequential Recommendation task, we follow the classical setting (Kang and McAuley, 2018; Sun et al., 2019) and select $100$ randomly sampled negative items for evaluation. We adopt two widely-used metrics Hit rate (HR@k) and Normalized Discounted Cumulative Gain (NDCG@k) with $k=5,10$.
@@ 4.3. Main Results
(5) Sequence modeling methods that possess analogous architectures and training objectives to those of behavior-tuned PLMs benefit more from our PLM-based initialization. Our item initialization is universal for different ID-based sequence modeling. Comparing the results based on SASRec and BERT4Rec, we find that our initialization facilitates more for BERT4Rec under the freeze setting. Besides, its FT (freezing) performs comparably to FT (trainable). We hypothesize that it is attributed to the similarities in sequence modeling architectures and training losses between RECFORMER and BERT4Rec. It also supports our hypothesis that PLMs may be excessively complex for behavior sequence modeling and could be substituted with simpler sequence models.
@@ 4.4. Results of Full-ranking Setting
The findings in Sec. 4.3 demonstrate the effectiveness of our proposed behavior-tuned PLMs initialized item embeddings under the random negative sampling setting. To further assess the robustness of our method, we conduct experiments with the full-ranking setting (Krichene and Rendle, 2020; Dallmann et al., 2021) (i.e., ranking the positive item with the whole item set) on the same datasets. From Table 4, we have:
@@
(2) Our method still consistently outperforms RECFORMER on all datasets. Compared with the random sampling setting in Table 3, the improvement in the full-ranking setting is not that impressive. This phenomenon is consistent with previous works (Li et al., 2023; Qu et al., 2023) that PLMs-based SR models perform well at making precise recommendations (i.e., pushing positive samples to the very top part of all candidates). It should be highlighted that our method greatly reduces the training and inference costs compared with RECFORMER. We only use PLM for item initialization, and the subsequent tuning and inference are only based on simplified sequence models without PLM. Considering our simple structure in serving (the same as SASRec), the current improvement is acceptable. The advantage of our method might be more activated in the matching stage (selecting the top hundreds of items from millions of candidates).
@@ 4.5. Analysis on Further Cooperations
Cooperating with trainable ID embeddings. Some previous works enhance item representation by adding the textual embeddings to its trainable item ID embeddings (Hou et al., 2022; Mu et al., 2023). Following this, we regard our behavior-tuned PLM initialization as the extra textual embeddings in SASRec and fix it during training (i.e., initialized embeddings as additional training signals). The results in Table 5 demonstrate that employing behavior-tuned PLMs initialized embeddings (i.e., PT and FT) leads to substantial performance enhancements, while embeddings from original PLMs (LF) do not. The results reconfirm that it is the behavior-aware semantics, not the original semantic information, that improves the SR performance.
@@
4.6. Analysis on Pre-training Sizes.
The Pre-training has been verified to benefit RECFORMER in SR on new domains. In Table 3, our PT setting also achieves satisfactory results. In this section, we explore the influence of pre-training dataset size on the effectiveness of our initialization by conducting experiments with PLMs pre-trained on randomly selected subsets of 0.01%, 0.1%, 1%, 2%, 10%, 20%, and 50% from the original pre-training dataset. From Fig. 2, we can observe that leveraging pre-training on behavioral datasets could significantly improve our performance. It is also astonishing that even a small amount of behavior-based pre-training (e.g., involving several thousand users) could successfully add the behavior-level preference to PLMs and achieve large improvement compared to the original Longformer. This observation reinforces the conclusion that the effectiveness of our method is primarily attributed to behavior-aware PLMs rather than the original PLMs. It also suggests that a minimal amount of pre-training behavioral data is possibly sufficient and transferable for new-domain SR tasks, implying that our method is practical and efficient based on a unified behavior-tuned PLM-based model.
@@ 5. CONCLUSION and FUTURE WORK
In this work, we deeply explore the effect of PLMs in SR, identifying significant model underutilization and redundancy of PLM in behavior sequence modeling by attention analyses and freezing/simplifying the sequence modeling of PLM. Next, we conduct extensive experiments and analyses on different PLM-based SR variants, discovering that the simplified sequence modeling of conventional ID-based SR models enhanced with the proposed behavior-tuned PLMs initialized item embeddings achieves a giant performance boost, while vanilla PLMs initializations are unable to obtain any improvements. In the future, we will explore the scalability of both SR datasets (including the size and domain diversity) and PLMs, determining whether any scaling laws apply to our proposed initialization. Additionally, identifying more appropriate training objectives to align PLMs with behavioral knowledge for initialization presents a promising avenue for research.
Acknowledgements.
This work is supported by the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001)